OpenStudy (anonymous):

Given the function f(x) = The quantity of 2x minus 5, divided by 3, which of the below expressions is correct?

OpenStudy (anonymous):

f−1(x) = The quantity of 5 minus 2x, divided by 3. f−1(x) = The quantity of 3x minus 5, divided by 2. f−1(x) = The quantity of 3x plus 5, divided by 2. f−1(x) = The quantity of negative 2x minus 5, divided by 3.

OpenStudy (marissalovescats):

Are you trying to find:  (inverse)

OpenStudy (anonymous):

yes

OpenStudy (marissalovescats):

Okay so how you solve for an inverse is you switch the place of the x and y (or x and f(x) because f(x) is technically y) and solve for f(x): And that's how you solve for the inverse.

OpenStudy (anonymous):

thank you very much

OpenStudy (marissalovescats):

You're very welcome